# Efficient first-principles inverse design of nanolasers

### v2

@@ 2.3 Nonlinear FOM via temporal coupled mode theory
To define a FOM that can be evaluated by solving a single linear system of equations and includes information about the coupling into the output channel, we use the expression for the amplitude in Eq. (10) in the context of TCMT. For a system with a sufficiently strong resonance (e.g. $Q\gtrsim 100$ [3]) coupled to a discrete set of input/output ports, TCMT derives a minimal set of equations coupling the resonant amplitude with the amplitudes of the port modes, constrained by energy conservation and reciprocity to be determined purely in terms of the coupling rates $1/\tau$ (lifetimes $\tau$[3]. Such a coupled optical system is illustrated in Fig. 2, where a cavity mode with amplitude $a_{0}$ (normalized so that $|a_{0}|^{2}$ is the energy in the cavity), total decay rate of the resonance $1/\tau_{0}$, and quality factor $Q=\omega_{0}\tau_{0}/2$, is coupled to a waveguide with an input amplitude $s_{\text{in}}$ and output amplitude $s_{\text{out}}$ (normalized so that $|s|^{2}$ is power) through a decay rate $1/\tau_{1}$. Note that the total decay rate 1/$\tau_{0}$ includes the decay rate $1/\tau_{1}$ into the waveguide plus the decay rates into other loss channels such as absorption and/or radiation [3].
Figure 2: Coupled optical system composed of a waveguide with input amplitude $s_{\text{in}}$ and output amplitude $s_{\text{out}}$, coupled to a cavity with an amplitude $a_{0}$ through the decay rate $\tau_{1}$.
@@ 2.3.1 Emission problem
In a direct formulation of the emission problem, the emitters in the gain medium will generate energy in the cavity mode, which then couples into the output waveguide mode. Using the TCMT equations, one can calculate the output power when there is no input waveguide mode ($s_{\text{in}}=0$), as [3]
$P_{\text{out}}=|s_{\text{out}}|^{2}=\frac{\omega_{0}}{Q_{1}}|a_{0}|^{2}\,,$ (11)
where $Q_{1}=\omega_{0}\tau_{1}/2$ is the quality factor for the coupling between cavity and waveguide. By substituting the expression for the amplitude in Eq. (10) and the expression for the threshold in Eq. (9), the TCMT output power becomes
$P_{\text{out}}\approx\Delta d\,\omega_{0}\frac{Q}{Q_{1}}\,\frac{\left(\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega\right)^{2}}{e_{c}^{-2}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{4}\,\mathop{}\!\mathrm{d}\Omega}.$ (12)
@@ 2.3.2 Reciprocal problem
Instead, we now consider the reciprocal problem, where the input port excites the resonance (often a useful approach for ensembles of emitters [15]). In this case, the TCMT equations imply a simple linear relationship between the input power and the resonant amplitude [3]
$U=|a_{0}|^{2}=\frac{2\tau_{0}^{2}}{\tau_{1}}|s_{\text{in}}|^{2}\ =\frac{1}{\omega_{0}}\frac{Q^{2}}{Q_{1}}P_{\text{in}}\,.$ (13)
In the high-$Q$ system, the field from the reciprocal solve $\mathbf{E}_{\mathrm{r}}$ is almost exactly equal to the cavity mode plus an $\mathcal{O}(1/\sqrt{Q})$ error [3]; thus, the energy $|a_{0}|^{2}$ can also be readily calculated by evaluating an integral of the reciprocal electric field in the cavity: $U=\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{r}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega$. Similarly, with this argument, the reciprocal field $\mathbf{E}_{\mathrm{r}}$ can also be used as a substitute for the modal field in Eqs. (912).
@@ 2.3.3 Nonlinear FOM for nanolaser optimization
where we have used the expression of the output power in Eq. (12), and the lasing threshold expression in Eq. (9). We have left out proportionality factors and have substituted the energy integral with the expression derived in Eq. (13). Although the “efficiency” ($P_{\text{out}}/P_{\text{pump}}$) on the left-hand-side of Eq. (14) is $\leq 1$, eliminating the proportionality factors results in the FOM defined by the right-hand-side taking arbitrary (nonnegative) values. Notice that all quality factors simplify, and we are left with an expression that only depends on reciprocal field integrals, meaning that the FOM can be evaluated with a single reciprocal solve (excited from the output waveguide). Moreover, this FOM $\sim|\mathbf{E}_{\mathrm{r}}|^{6}/|\mathbf{E}_{\mathrm{r}}|^{4}\sim|\mathbf{E}_{\mathrm{r}}|^{2}$ is roughly proportional to the energy in the cavity ($\sim|\mathbf{E}_{\mathrm{r}}|^{2}$) and thus to the $Q$ factor. This fact will ensure that maximizing the FOM yields high-$Q$ optimized cavities, making the perturbative single-pole approximation more and more accurate as the optimization progresses.
@@ 2.3.4 LDOS and the single-emitter limit
Interestingly, in the single-emitter limit, where the gain distribution is modeled as being concentrated at a single point $\mathbf{r^{\prime}}$, i.e. $D_{0}(\mathbf{r})\propto\delta(\mathbf{r}-\mathbf{r^{\prime}})$, the nonlinear FOM (Eq. (14)) and the naive FOM (Eq. (15)) become equivalent, reducing to an LDOS-like FOM when solving the reciprocal problem [45]: $\text{FOM}\propto|\mathbf{E}_{\mathrm{r}}(\mathbf{r^{\prime}})|^{2}$. Whereas the usual definition of LDOS would correspond to the total power expended by an orientation-averaged dipole source at $\mathbf{r^{\prime}}$ [5], this FOM corresponds (via reciprocity [45, App. C]) to the dipole power emitted only into the output channel, a form of extraction efficiency. Such a point-gain FOM is therefore reminiscent of earlier efforts [20, 21, 22, 23, 24] to optimize cavities by maximizing LDOS or extraction efficiency.
@@ 2.4 Accounting for steady-state diffusion in the gain medium
In semiconductor lasers with extended media, i.e., as opposed to isolated emitters such as quantum dots, it is paramount to account for diffusion of excited carriers. Thus, while the spatial profile of the lasing mode leads to a spatial variation of the local rate of stimulated emission and hence a spatially varying depletion of excited states, gain diffusion smoothens out the effective gain profile. As introduced in the complex-SALT (C-SALT) model [4], this effect can be accounted for by modeling a bulk semiconductor gain medium in the free-carrier approximation, where carrier–carrier Coulomb interactions are neglected. In this approximation, one can include diffusion in Eq. (7) by applying a set of linear operators
@@
Similar to the derivations in Sec. 2.2, we calculate the pumping strength needed to reach the lasing threshold by employing perturbation theory for small $|\mathbf{E}_{L}|^{2}\sim|a_{0}|^{2}$
$d_{\text{thresh}}=\frac{1}{Q}\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}\mathbb{S}[D_{0}](\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,.$ (17)
@@ 2.5 Topology optimization formulation
In this section, we review the standard density-based TopOpt formulation [46, 26] that we employ to describe and optimize our degrees of freedom, which is mostly independent of the choice of FOM. Density-based TopOpt parameterizes the distribution of materials in space by introducing a design field, the “density” $\rho(\mathbf{r})\in[0,1]$, where the extremes $\rho=0$ and $\rho=1$ correspond to the two materials, while unphysical intermediate materials are also permitted temporarily as the design evolves. To regularize the optimization problem by introducing a weak sense of geometric lengthscale (ensuring convergence as the computational resolution is increased), one applies a filtering and thresholding scheme to the design variables. The thresholding, which pushes the design towards a “binarized” structure in which intermediate material are excluded, is performed via a smoothed Heaviside projection [47]
@@
where $\hat{\rho}$ is the physical field, $\Theta_{\beta,\eta}(\tilde{\rho})$ is the thresholding function, $\beta\in[1,\infty)$ and $\eta\in[0,1]$ are hyper-parameters that control the threshold sharpness and value respectively. The filtered design field $\tilde{\rho}$, which regularizes the problem by introducing a minimum lengthscale $r_{\text{f}}$, is obtained using a Helmholtz-based filter [48]
$-\left(\frac{r_{\text{f}}}{2\sqrt{3}}\right)^{2}\nabla^{2}\tilde{\rho}+\tilde{\rho}=\rho\,,$ (20)
where $r_{\text{f}}$ is a filter radius. The projection steepness $\beta$ is gradually increased as the optimization progresses: a small $\beta$ allows grayscale structures in which the topology can change smoothly, while a larger $\beta$ binarizes the structure but slows convergence (so the largest $\beta$ is only imposed at the end, when the structure is nearly converged) [49, 50]; the precise optimization details are given in App. D.
The filtering step regularizes the problem, but does not strictly exclude features that have arbitrarily small lengthscales. In order to exclude such non-manufacturable results, we employ a standard technique to impose additional geometric minimum-lengthscale constraints  [51, 52] with a minimum feature size just above the single-pixel level ($\gtrsim 40$ nm) on both the solid and void regions, with the details in App. D.
@@
where subscript $1$ denotes the material and 00 the background. We also model a design-dependent extinction coefficient
$\kappa=-(\alpha+\alpha^{\prime}\hat{\rho}(1-\hat{\rho}))\,,$ (22)
where $\alpha$ is a coefficient that introduces artificial optical losses to broaden the frequency response of high-$Q$ optical resonances, thus facilitating the optimization problem [21]; and $\alpha^{\prime}$ is an attenuation factor that discourages intermediate values of $\hat{\rho}$ by introducing optical losses. As the optimization progresses, $\alpha$ is gradually “turned off” to zero (eliminating the unphysical damping) while $\alpha^{\prime}$ is gradually “turned on” (penalizing unphysical intermediate materials). The gain medium is also dependent on $\hat{\rho}$, because we assume that gain is only present in the solid material. Furthermore, in Sec. 3, we explore gain regions of varying diameters, described by a Gaussian profile around the center $\mathbf{r}_{0}$ of the cavity, yielding the equation
$D_{0}(\mathbf{r},\hat{\rho})=\varepsilon_{\mathrm{r}}(\mathbf{r})\hat{\rho}(\mathbf{r})\,e^{-|\mathbf{r}-\mathbf{r}_{0}|^{2}/2\sigma_{\text{g}}^{2}}\,,$ (23)
@@ 3 Computational Results
In this section, we apply the theoretical framework reviewed and developed in the previous sections to optimize nanocavity-based laser devices. First, we compare our developed first-principles FOM with a heuristic generalization of the LDOS (Sec. 2.3.4) through the inverse design 2d cavities (Sec. 3.1), demonstrating the increasing importance of a first-principles model as the diameter of the gain region increases. Then, we study how steady-state diffusion effects influence the optimized nanocavity designs (Sec. 3.2). In all cases, the permittivity of the solid regions is $\varepsilon=12$ (characteristic of semiconductors) and of the void regions is $\varepsilon=1$ (air). The gain medium material is modeled as a III–V semiconductor-like material with a direct bandgap. Lastly (Sec. 3.3), we consider the inverse design of more realistic 3d cavities in semiconductor-on-insulator system (SOI, $\varepsilon=1.44^{2}$ for the substrate). (Detailed simulation and optimization parameters are given in App. D.)
In all cases, the laser is optimized for a wavelength of $1550$ nm, and we impose a minimum lengthscale [51] (Sec. 2.5) just above the single-pixel level ($\gtrsim 40$ nm) on both solid and void regions. The design region is $3.1\,\upmu\mathrm{m}\times 3.1\,\upmu\mathrm{m}$, connected to a single-mode output waveguide (width 500 nm), with PML absorbing boundary layers [53, 54] along with absorbing boundary conditions [55]. The 2d simulations employ the $H_{z}$ (out-of-plane $\mathbf{H}$, in-plane $\mathbf{E}$) polarization, whereas the 3d simulations use the analogous mostly-$H_{z}$ waveguide mode. All simulations employ a finite-element discretization whose details are given in App. D, with resolution $\approx 40$ nm in the design region. Antisymmetric mirror-symmetry boundary conditions (Neumann for $H_{z}$, equivalent to a perfect electric conductor) are imposed to halve the computational region (bisecting the waveguide). Although such a non-convex optimization problem may, in principle, have non-symmetric local optima (and TopOpt often exhibits many local optima with similar performance [56, 57]), we found that symmetric optima were often found even if we did not impose symmetry, and imposing mirror symmetry reduces computational costs. Except where otherwise noted, all optimizations began with a homogeneous $\rho=0.5$ starting structure.
@@
As pointed out in Sec. 2.3.4, in the single-emitter limit where the active region is point-like, the naive and nonlinear FOM become equivalent, and in consequence we observe in Fig. 3(b) that the optimized cavities become nearly equivalent for small $\sigma_{\text{g}}\ll\lambda$. In this single-emitter limit, the cavity geometry exhibits a central bowtie-like feature—a structure commonly arising in LDOS optimization [20, 21, 22, 23, 24]—which is known to enhance light–-matter interaction at a single point [58, 59, 30] by concentrating the in-plane electric field at the emitter location thanks to a field singularity [60, 9] at any sharp tip. (Here, the bowtie sharpness is limited by our $\gtrsim 40$ nm lengthscale constraints.) Due to the strong field concentration at the cavity-tip, the normalized FOM/$\zeta(\sigma_{g})$, has a larger value compared to the cavities that were optimized for larger gain-region sizes.
@@
As the active region size increases, the system transitions beyond the single-emitter regime, and the cavity designs optimized for the naive and nonlinear FOMs begin to differ significantly as the gain diameter approaches the wavelength ($\sigma_{\text{g}}\gtrsim\lambda/n_{1}\sim 450$ nm). In both cases, the spatial averaging of the emission means that arbitrarily sharp tips are no longer beneficial (because $\int|\mathbf{E}|^{2}$ is finite) [30]. For our largest gain diameter [$\sigma_{\text{g}}=500$ nm, Fig. 3(d)], the structure optimized using the nonlinear FOM achieves a performance 3$\approx$ 3$\times$ 3 × greater than the structure optimized with the naive FOM. We attribute this to the nonlinear “hole-burning” effect (Sec. 2.1) captured in the denominator of the nonlinear FOM (Eq. (14)) but not present in the naive FOM (Eq. (15)), which penalizes strong spatial field localization. As illustrated in Figure 4, the electric-field distribution shows that the cavity optimized for the naive FOM concentrates the field at its center, whereas the cavity optimized for the nonlinear FOM distributes the field more homogeneously around the active region $D_{0}$, reducing hole burning and leading to a higher nonlinear FOM. Moreover, as the active region size becomes larger, the normalized FOM ($\text{FOM}/\zeta(\sigma_{g})$) decreases [Fig. 3(a)], and transitions from the sharp-tip large-enhancement regime to distributing the electric field within the active material, which results in the normalized FOM becoming roughly proportional to $Q$ for larger active region sizes (Sec. 2.3.3). Note that the non-normalized FOM increases by several orders of magnitude as the active‐region volume grows, simply reflecting the larger gain medium emitting light into the lasing mode.
@@
Finally, we validate a posteriori one of the main assumptions in the first-principles modeling based on the SPA-SALT (Sec. 2.2): that the optimized cavities are indeed high-$Q$ cavities. By means of a post-processing eigensolve on the optimized the cavities, we find that the $Q$s of the lasing modes are in the range $Q\in[350,1000]\gtrsim 100$, which validates the SPA-SALT and TCMT approximations [42, 3]. Not only that, as pointed out in Sec. 1, there is a trade-off between designing a high-$Q$ cavity and a cavity that couples efficiently to an output port. By computing the power fraction in the waveguide relative to the total lasing-mode power in the post-processing eigensolve, we evaluate the lasing mode’s optical extraction efficiency in the topology-optimized cavities. As expected from the definition of the nonlinear FOM (Eq. (14)), which directly incorporates out-coupling effects, we find values of the optical extraction efficiency $\geq 0.9$ for most of the designs, highlighting the ability of the FOM to yield high-$Q$ cavities with high optical extraction efficiencies.
@@
Gain diffusion is important in extended gain media, such as bulk materials and quantum wells, when the extent of the gain region becomes comparable to or larger than the diffusion length. In such cases, the diffusion counteracts the effect of spatial-hole burning. To understand the impact of gain diffusion on the optimization results we use a diffusion model based on the C-SALT framework [4](Sec. 2.4) to analyze the performance of the cavity previously optimized for a gain medium of size $\sigma_{\text{g}}=250$ nm (Fig. 3), where diffusion was neglected. Using this design as an initial guess, we re-optimize the cavity in Fig. 5 using the diffusion FOM (Eq. (18)), assuming a diffusion length of $R_{0}=5\,\upmu\text{m}$ comparable to typical electron and hole diffusion lengths in semiconductor materials such as Si [61, 62], InP [62], (In,Ga)N [63], InAs [62], GaAs [62], and InGaAs [64]. Note that as the diffusion lengths decrease, the system approaches the non-diffusive limit ($R^{2}_{\nabla}\ll 1$, Eq. (16)), and the optimization results will converge toward those presented in Sec. 3.1.

### v3

@@ 2.3 Nonlinear FOM via temporal coupled mode theory
[…]
Figure 2: Coupled optical system composed of a waveguide guiding an input field with amplitude $s_{\text{in}}$ and an output field with amplitude $s_{\text{out}}$, coupled to a cavity with an amplitude $a_{1}$ through the decay rate $\tau_{\text{wg}}$.
@@ 2.3.1 Emission problem
In a direct formulation of the emission problem, the emitters in the gain medium will generate energy in the cavity mode, which then couples into the output waveguide mode. Using the TCMT equations, one can calculate the output power when there is no input waveguide mode ($s_{\text{in}}=0$), as [phot_crys]
$P_{\text{out}}=|s_{\text{out}}|^{2}=\frac{\omega_{1}}{Q_{\text{wg}}}|a_{1}|^{2}\,$, (11)
where $Q_{\text{wg}}=\omega_{1}\tau_{\text{wg}}/2$ is the quality factor for the coupling between cavity and waveguide. By substituting the expression for the amplitude in Eq. (10) and the expression for the threshold in Eq. (9), the TCMT output power becomes
$P_{\text{out}}\approx\Delta d\,\omega_{1}\frac{Q}{Q_{\text{wg}}}\,\frac{\left(\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega\right)^{2}}{e_{c}^{-2}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{4}\,\mathop{}\!\mathrm{d}\Omega}.$ (12)
@@ 2.3.2 Reciprocal problem
Instead, we now consider the reciprocal problem, where the input port excites the resonance (often a useful approach for ensembles of emitters [Yao_2022]). In this case, the TCMT equations imply a simple linear relationship between the input power and the resonant amplitude [phot_crys]
$U=|a_{1}|^{2}=\frac{2\tau_{1}^{2}}{\tau_{\text{wg}}}|s_{\text{in}}|^{2}\ =\frac{1}{\omega_{1}}\frac{Q^{2}}{Q_{\text{wg}}}P_{\text{in}}\,.$ (13)
In the high-$Q$ system, the field from the reciprocal solve $\mathbf{E}_{\mathrm{r}}$ is almost exactly equal to the cavity mode plus an $\mathcal{O}(1/\sqrt{Q})$ error [phot_crys]; thus, the energy $|a_{1}|^{2}$ can also be readily calculated by evaluating an integral of the reciprocal electric field in the cavity: $U=\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{r}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega$. Similarly, with this argument, the reciprocal field $\mathbf{E}_{\mathrm{r}}$ can also be used as a substitute for the modal field in Eqs. (912).
@@ 2.3.3 Nonlinear FOM for nanolaser optimization
where we have used the expression of the output power in Eq. (12), and the lasing threshold expression in Eq. (9). We have left out proportionality factors and have substituted the energy integral with the expression derived in Eq. (13). Although the “efficiency” ($P_{\text{out}}/P_{\text{pump}}$) on the left-hand-side of Eq. (14) is $\leq 1$, eliminating the proportionality factors results in the FOM defined by the right-hand-side taking arbitrary (nonnegative) values. Notice that all quality factors simplify, and we are left with an expression that only depends on reciprocal field integrals, meaning that the FOM can be evaluated with a single reciprocal solve (excited from the output waveguide). Moreover, this FOM $\sim|\mathbf{E}_{\mathrm{r}}|^{6}/|\mathbf{E}_{\mathrm{r}}|^{4}\sim|\mathbf{E}_{\mathrm{r}}|^{2}$ is roughly proportional to the energy in the cavity ($\sim|\mathbf{E}_{\mathrm{r}}|^{2}$) and thus to the $Q$ factor. This fact will ensure that maximizing the FOM yields high-$Q$ optimized cavities, making the perturbative single-pole approximation more and more accurate as the optimization progresses (numerical validation of this high-$Q$ approximation is provided in App. D for optimized nanolasers).
@@ 2.3.4 LDOS and the single-emitter limit
Interestingly, in the single-emitter limit, where the gain distribution is modeled as being concentrated at a single point $\mathbf{r^{\prime}}$, i.e. $D_{0}(\mathbf{r})\propto\delta(\mathbf{r}-\mathbf{r^{\prime}})$, the nonlinear FOM (Eq. (14)) and the naive FOM (Eq. (15)) become equivalent, reducing to an LDOS-like FOM when solving the reciprocal problem [LDOS_reciprocal]: $\text{FOM}\propto|\mathbf{E}_{\mathrm{r}}(\mathbf{r^{\prime}})|^{2}$. Whereas the usual definition of LDOS would correspond to the total power expended by an orientation-averaged dipole source at $\mathbf{r^{\prime}}$ [OskooiJo13-sources], this FOM corresponds (via reciprocity [LDOS_reciprocal, App. C]) to the dipole power emitted only into the output channel, a form of extraction efficiency. Such a point-gain FOM is therefore reminiscent of earlier efforts [LDOS_opt_wang, LDOS_opt_liang, LDOS_bounds, albrechtsen_nanometer-scale_2022, LDOS_G] to optimize cavities by maximizing LDOS or extraction efficiency.
@@ 2.4 Accounting for steady-state diffusion in the gain medium
In semiconductor lasers with extended media, i.e., as opposed to isolated emitters such as quantum dots, it is paramount to account for diffusion of excited carriers. Thus, while the spatial profile of the lasing mode leads to a spatial variation of the local rate of stimulated emission and hence a spatially varying depletion of excited states, gain diffusion smoothens out the effective gain profile. As introduced in the complex-SALT (C-SALT) model [csalt], this effect can be accounted for by modeling a bulk semiconductor gain medium in the free-carrier approximation, where carrier–carrier Coulomb interactions are neglected. In this approximation, one can include diffusion in Eq. (7) by applying a set of linear operators
@@
Similar to the derivations in Sec. 2.2, we calculate the pumping strength needed to reach the lasing threshold by employing perturbation theory for small $|\mathbf{E}_{L}|^{2}\sim|a_{1}|^{2}$
$d_{\text{thresh}}=\frac{1}{Q}\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}\mathbb{S}[D_{0}](\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,.$ (17)
@@ 2.5 Topology optimization formulation
In this section, we review the standard density-based TopOpt formulation [topopt_dens, jensen_topology_2011] that we employ to describe and optimize our degrees of freedom, which is mostly independent of the choice of FOM. Density-based TopOpt parameterizes the distribution of materials in space by introducing a design field, the “density” $\rho(\mathbf{r})\in[0,1]$, where the extremes $\rho=0$ and $\rho=1$ correspond to the two materials, while unphysical intermediate materials are also permitted temporarily as the design evolves. To regularize the optimization problem by introducing a weak sense of geometric lengthscale (ensuring convergence as the computational resolution is increased), one applies a filtering and thresholding scheme to the design variables. The thresholding, which pushes the design towards a “binarized” structure in which intermediate material are excluded, is performed via a smoothed Heaviside projection [projection]
@@
[…]
$\left(\frac{r_{\text{f}}}{2\sqrt{3}}\right)^{2}\nabla^{2}\tilde{\rho}+\tilde{\rho}=\rho\,$, (20)
where $r_{\text{f}}$ is a filter radius. The projection steepness $\beta$ is gradually increased as the optimization progresses: a small $\beta$ allows grayscale structures in which the topology can change smoothly, while a larger $\beta$ binarizes the structure but slows convergence (so the largest $\beta$ is only imposed at the end, when the structure is nearly converged) [ole_beta, HammondOs21]; the precise optimization details are given in App. E.
The filtering step regularizes the problem, but does not strictly exclude features that have arbitrarily small lengthscales. In order to exclude such non-manufacturable results, we employ a standard technique to impose additional geometric minimum-lengthscale constraints  [lengthscale, meep] with a minimum feature size just above the single-pixel level ($\gtrsim 40$ nm) on both the solid and void regions, with the details in App. E.
@@
where subscript $1$ denotes the material and 0 the background. We also model a design-dependent extinction coefficient
$),\kappa=-(\alpha+\alpha^{\prime}\hat{\rho}(1-\hat{\rho}$))\,, (22)
where $\alpha$ is a coefficient that introduces artificial optical losses to broaden the frequency response of high-$Q$ optical resonances, thus facilitating the optimization problem [LDOS_opt_liang]; and $\alpha^{\prime}$ is an attenuation factor that discourages intermediate values of $\hat{\rho}$ by introducing optical losses. As the optimization progresses, $\alpha$ is gradually “turned off” to zero (eliminating the unphysical damping) while $\alpha^{\prime}$ is gradually “turned on” (penalizing unphysical intermediate materials). The gain medium is also dependent on $\hat{\rho}$, because we assume that gain is only present in the solid material. Furthermore, in Sec. 3, we explore gain regions of varying diameters, described by a Gaussian profile around the center $\mathbf{r}_{0}$ of the cavity, yielding the equation
$D_{0}(\mathbf{r},\hat{\rho})=\varepsilon_{\mathrm{c}(\mathbf{r})}\hat{\rho}(\mathbf{r})\,e^{-|\mathbf{r}-\mathbf{r}_{0}|^{2}/2\sigma_{\text{g}}^{2}}\,$, (23)
@@ 3 Computational Results
In this section, we apply the theoretical framework reviewed and developed in the previous sections to optimize nanocavity-based laser devices. First, we compare our developed first-principles FOM with a heuristic generalization of the LDOS (Sec. 2.3.4) through the inverse design 2d cavities (Sec. 3.1), demonstrating the increasing importance of a first-principles model as the diameter of the gain region increases. Then, we study how steady-state diffusion effects influence the optimized nanocavity designs (Sec. 3.2). In all cases, the permittivity of the solid regions is $\varepsilon=12$ (characteristic of semiconductors) and of the void regions is $\varepsilon=1$ (air). The gain medium material is modeled as a III–V semiconductor-like material with a direct bandgap. Lastly (Sec. 3.3), we consider the inverse design of more realistic 3d cavities in semiconductor-on-insulator system (SOI, $\varepsilon=1.44^{2}$ for the substrate). (Detailed simulation and optimization parameters are given in App. E.)
In all cases, the laser is optimized for a wavelength of $1550$ nm, and we impose a minimum lengthscale [lengthscale] (Sec. 2.5) just above the single-pixel level ($\gtrsim 40$ nm) on both solid and void regions. The design region is $2.325\,\upmu\mathrm{m}\times 2.325\,\upmu\mathrm{m}$, connected to a single-mode output waveguide (width 500 nm), with PML absorbing boundary layers [taflove, BERENGER1994185] along with absorbing boundary conditions [jin]. The 2d simulations employ the $H_{z}$ (out-of-plane $\mathbf{H}$, in-plane $\mathbf{E}$) polarization, whereas the 3d simulations use the analogous mostly-$H_{z}$ waveguide mode. All simulations employ a finite-element discretization whose details are given in App. E, with resolution $\approx 40$ nm in the design region. Antisymmetric mirror-symmetry boundary conditions (Neumann for $H_{z}$, equivalent to a perfect electric conductor) are imposed to halve the computational region (bisecting the waveguide). Although such a non-convex optimization problem may, in principle, have non-symmetric local optima (and TopOpt often exhibits many local optima with similar performance [ChenCh24, dalklint2024performanceboundstopologyoptimization]), we found that symmetric optima were often found even if we did not impose symmetry, and imposing mirror symmetry reduces computational costs. Except where otherwise noted, all optimizations began with a homogeneous $\rho=0.5$ starting structure.
@@
To showcase the difference of optimizing for a naive generalization of the LDOS (Eq. (15), Sec. 2.3.4) or for the first-principles nonlinear FOM (Eq. (14), Sec. 2.3.3), in Fig. 3 we compare the performance of optimized nanolasers for different gain region sizes: $\sigma_{\text{g}}\in\{25\text{ nm},50\text{ nm},100\text{ nm},250\text{ nm},400\text{ nm}$, $500\text{ nm}\}$. Half of the cavities are optimized for the naive FOM (green) , while the other half are optimized for the derived nonlinear FOM (yellow). To make the FOM easier to interpret, reproduce, and compare across nanolaser platforms, we normalize the nonlinear FOM by dividing by $\zeta(\sigma_{g})=\max\left\{|\mathbf{E}_{\text{in}}|^{2}\right\}\left(\int_{\Omega}D^{*}_{0}(\sigma_{g})\mathop{}\!\mathrm{d}\Omega\right)^{2}$, where $\mathbf{E}_{\mathrm{in}}$ is the input‐waveguide electric field and $D^{*}_{0}=e^{-|\mathbf{r}-\mathbf{r}_{0}|^{2}/2\sigma_{\text{g}}^{2}}$ is the Gaussian gain profile with the structure ($\varepsilon_{\mathrm{r}}\hat{\rho}$) dependence removed. Note that the results have been cross checked to ensure that indeed a design optimized for a certain domain size is the best for that domain size.
@@
As the active region size increases, the system transitions beyond the single-emitter regime, and the cavity designs optimized for the naive and nonlinear FOMs begin to differ significantly as the gain diameter approaches the wavelength ($\sigma_{\text{g}}\gtrsim\lambda/n_{1}\sim 450$ nm). In both cases, the spatial averaging of the emission means that arbitrarily sharp tips are no longer beneficial (because $\int|\mathbf{E}|^{2}$ is finite) [YaoVe23]. For our largest gain diameter $\sigma_{\text{g}}=500$ nm, Fig. 3(d)], the structure optimized using the nonlinear FOM achieves a performance $\approx 3\times$ greater than the structure optimized with the naive FOM. We attribute this to the nonlinear “hole-burning” effect (Sec. 2.1) captured in the denominator of the nonlinear FOM (Eq. (14)) but not present in the naive FOM (Eq. (15)), which penalizes strong spatial field localization. As illustrated in Figure 4, the electric-field distribution shows that the cavity optimized for the naive FOM concentrates the field at its center, whereas the cavity optimized for the nonlinear FOM distributes the field more homogeneously around the active region $D_{0}$, reducing hole burning and leading to a higher nonlinear FOM. Moreover, as the active region size becomes larger, the normalized FOM ($\text{FOM}/\zeta(\sigma_{g})$) decreases [Fig. 3(a)], and transitions from the sharp-tip large-enhancement regime to distributing the electric field within the active material, which results in the normalized FOM becoming roughly proportional to $Q$ for larger active region sizes (Sec. 2.3.3). Note that the non-normalized FOM increases by several orders of magnitude as the active‐region volume grows, simply reflecting the larger gain medium emitting light into the lasing mode.
@@
Finally, we validate a posteriori one of the main assumptions in the first-principles modeling based on the SPA-SALT (Sec. 2.2): that the optimized cavities are indeed high-$Q$ cavities. By means of a post-processing eigensolve on the optimized the cavities, we find that the $Q$s of the lasing modes are in the range $Q\in[350,1000]\gtrsim 100$, which validates the SPA-SALT and TCMT approximations [cerjan_2016, phot_crys] as shown in App. D. Not only that, as pointed out in Sec. 1, there is a trade-off between designing a high-$Q$ cavity and a cavity that couples efficiently to an output port. By computing the power fraction in the waveguide relative to the total lasing-mode power in the post-processing eigensolve, we evaluate the lasing mode’s optical extraction efficiency in the topology-optimized cavities. As expected from the definition of the nonlinear FOM (Eq. (14)), which directly incorporates out-coupling effects, we find values of the optical extraction efficiency $\geq 0.9$ for most of the designs, highlighting the ability of the FOM to yield high-$Q$ cavities with high optical extraction efficiencies. We finally note that we have also assessed fabrication robustness by evaluating how systematic over– and under–etching would affect performance using the nonlinear FOM for the device with $\sigma=500$ nm. Following [eta], a change in threshold value $\Delta\eta=0.1$ corresponds to a dimensional variation of $\Delta l\approx 10$ nm, yielding a FOM scaled by a factor $\approx\times 0.85$ per nanometer. This provides a quantitative estimate of expected performance tolerance to nanoscale variations for the optimized high-$Q$ nanolaser designs.
@@
Gain diffusion is important in extended gain media, such as bulk materials and quantum wells, when the extent of the gain region becomes comparable to or larger than the diffusion length. In such cases, the diffusion counteracts the effect of spatial-hole burning. To understand the impact of gain diffusion on the optimization results we use a diffusion model based on the C-SALT framework [csalt](Sec. 2.4) to analyze the performance of the cavity previously optimized for a gain medium of size $\sigma_{\text{g}}=250$ nm (Fig. 3), where diffusion was neglected. Using this design as an initial guess, we re-optimize the cavity in Fig. 5 using the diffusion FOM (Eq. (18)), assuming a diffusion length of $R_{0}=5\,\upmu\text{m}$ comparable to typical electron and hole diffusion lengths in semiconductor materials such as Si [Diff_Si, semicon_params], InP [semicon_params], (In,Ga)N [PhysRevApplied.15.054015], InAs [semicon_params], GaAs [semicon_params], and InGaAs [PhysRevB.70.205311]. Note that as the diffusion lengths decrease, the system approaches the non-diffusive limit ($R^{2}_{\nabla}\ll 1$, Eq. (16)), and the optimization results will converge toward those presented in Sec. 3.1.